# NeuronMM: High-Performance Matrix Multiplication for LLM Inference on AWS Trainium

### v2

@@ 1. Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of text-based tasks (liu2024deepseek, ; yang2025qwen3, ). Yet, the steady growth in their parameter counts and architectural complexity makes deployment increasingly prohibitive, particularly in resource-constrained environments. These challenges have driven extensive research into model compression (agarwal2024policy, ; frantar2023sparsegpt, ; dettmers2022gpt3, ; dettmers20218, ; sanh2020movement, ; zhang2020accelerating, ; li2023symbolic, ; chen2022disco, ) and efficient hardware design (wan2023efficient, ; zhu2024survey, ; jouppi2020domain, ; bai2023transformer, ; tzanos2022hardware, ; tuli2023acceltran, ; liu2023hardsea, ; ji2024beta, ; tuli2023acceltran, ).
AI accelerators (jouppi2023tpu, ; coburn2025meta, ; firoozshahian2023mtia, ), customized to AI workloads, provide cost-effective and high-performance solutions for training and inference. Trainium is an AI accelerator recently developed by Amazon Web Services (AWS). It has been reported that Trainium can deliver 30–40% lower cost while providing performance comparable to GPU-based EC2 instances (aws_trainium, ). Each Trainium2 device (the most advanced Trainium) integrates two NeuronCores, each delivering up to 95 TFLOPS of FP16/BF16 compute capability, comparable to NVIDIA A100 GPUs at roughly 60% of the cost. Such a cost–effective advantage makes Trainium an attractive platform for LLM training and inference (fu2024distributed, ; fan2024hlat, ). Furthermore, Trainium, as a typical systolic-array architecture, features a programmable memory hierarchy (including two types of on-chip SRAMs and off-chip HBM). It also provides a rich set of specialized compute engines tailored for various AI operators. Such hardware heterogeneity gives programmers a lot of flexibility to explore for better performance.
@@ 2.1. SVD for Weight Compression
SVD is a well-established technique for approximating high-dimensional matrices with low-rank representations (demmel1997applied, ). Given a weight matrix $W$, SVD factorizes it into three matrices: $U$, $\Sigma$, and $V$, such that $W=U\Sigma V^{\top}$. By retaining only the top-$k$ singular values in $\Sigma$ and their corresponding singular vectors in $U$ and $V$, one obtains a low-rank approximation $W\approx U_{k}\Sigma_{k}V_{k}^{\top}$. This approximation preserves the most informative components of $W$ while substantially reducing the number of parameters to represent $W$. As a result, SVD is particularly well-suited for compressing the large weight matrices in the linear layers of LLMs, where parameter reduction directly improves efficiency with tolerable accuracy loss.
@@ 2.2. AWS Trainium
AWS Trainium is a custom silicon chip designed to accelerate deep learning workloads. It adopts a systolic array–based architecture with rich hardware heterogeneity. Each Trainium chip integrates two NeuronCores. Each NeuronCore functions as an independent heterogeneous compute unit composed of a rich set of specialized engines designed for different operations, such as tensor engine, scalar engine, vector engine, and GPSIMD engine. In addition, Trainium includes DMA engines that can transfer data between HBM and on-chip SRAM in Trainium. Those engines operate in parallel, enabling Trainium to efficiently support diverse deep learning tasks (aws_trainium_arch, ). In addition, we particularly focus on the tensor engine for matmul in the following discussion.
We use Neuron Kernel Interface (NKI) (aws2024neuron, ), a bare-metal language and compiler for directly programming NeuronDevices available on AWS Trainium, in our study.
@@
Each tile-level matmul involves two input matrices, named as the stationary (left-hand side) matrix and moving (right-hand side) matrix (using AWS terms). The tensor engine loads the stationary matrix into its internal storage and streams the moving matrix across it. Due to the hardware’s systolic array design, the stationary matrix must be consumed in a transposed layout (aws_trainium_arch, ). For clarity, we denote the low-level instruction nki.isa.nc_matmul (aws_nki_matrix_multiplication, ) as NKIMatmul, which carries out this tile-level operation. Thus, computing the product of $AB$ is formulated as NKIMatmul(stationary = $A^{T}$, moving = $B$). The stationary matrix, which remains fixed during the computation, is transposed so that its rows align with the columns of the moving matrix, ensuring both input tiles share the same first dimension which corresponds to the partition dimension in SBUF.
@@
The tile is the fundamental, hardware-native unit of data processed by a single NKIMatmul instruction. The tensor engine imposes strict, hardware-defined maximum dimensions on these tiles, and these constraints differ for the two matrix operands (aws_trainium_arch, ). Specifically, a tile for the stationary matrix cannot exceed dimensions of (128, 128) for its partition and free axes, respectively. In contrast, a tile for the moving matrix can have a larger free axis, with its dimensions limited to (128, 512). When a matrix is larger than these hardware limits, it must be divided into multiple tiles for processing.
@@ 3.1. Challenge 1: I/O Bottleneck
Executing the SVD-compressed $XUV$ computation with a standard, sequential approach on Trainium results in poor hardware utilization. Figure 4 shows a Neuron Profiler trace of the SVD-compressed up_projection matmul in Deepseek-V3 (liu2024deepseek, ). For an input length of 4096, with hidden size 7168 and intermediate size 18432, the up_projection matrix ([7168, 18432]) is factorized into two low-rank matrices: $U$ ([7168, 4096]) and $V$ ([4096, 18432]). This transforms the original matmul ([4096, 7168] $\times$ [7168, 18432]) into a chain of three multiplications: [4096, 7168] $\times$ [7168, 4096] $\times$ [4096, 18432]. As highlighted in Figure 4, this decomposition results in frequent idle gaps in Model FLOPs Utilization (MFU) and inefficient use of the on-chip SBUF, indicating that the tensor engine is often stalled waiting for data.
@@ 4. Design
NeuronMM compresses the large weight matrices in MLP layers using block-aligned SVD and restores accuracy through Low-Rank Adaptation (LoRA) (hu2022lora, ) fine-tuning. It further provides high-performance NKI kernels, termed TrainiumFusion, that execute matmuls on compressed models efficiently by exploiting Trainium’s architectural features.
@@ 4.1. Block-Aligned SVD
Leveraging existing work. Following the standard workflow of post-training LLM compression methods (yuan2023asvd, ; wang2024svd, ; ding2025dipsvd, ), the weight matrix $W$ is first scaled by a matrix $S$ to capture the influence of input activations. $S$ is derived from a random set of input sentences. Specifically, for each MLP layer in LLMs, we record input activations $X$ using forward hooks, compute the covariance matrix, and apply Cholesky decomposition:
@@
where $\text{Model}\_\text{size}_{\text{compressed}}$ and $\text{Model}\_\text{size}_{\text{original}}$ denote the model parameter counts with and without compression. This definition follows the existing work on LLM compression (wang2024svd, ; wang2025svd, ; li2025adasvd, ; yuan2023asvd, ; ding2025dipsvd, ).
@@
Although SVD substantially reduces the size of the weight matrix, it inevitably introduces accuracy loss. To recover accuracy, we apply LoRA fine-tuning to the compressed weights, similar to the prior work (wang2024svd, ; li2025adasvd, ). During fine-tuning, we freeze the compressed weights $U_{r}^{\prime}$ and $V_{r}^{\prime}$, and fine-tune them with LoRA:
@@ 4.2.1. XUV NKI Kernel
Using Equations 10-11, we determine the optimal block size. By the roofline model, a kernel becomes compute-bound when its arithmetic intensity exceeds a hardware-specific threshold (williams2009roofline, )—for a Trainium NeuronCore with bfloat16 data, this threshold is 222 Flops/Byte (nki_mm2025, ). Setting the arithmetic intensity (Equation 10) to this threshold yields the minimum block size $B_{M}$ required to saturate the tensor engine. Starting with this $B_{M}$, the peak SBUF usage model (Equation 11) can then identify block combinations that maximize data reuse within the 24 MB SBUF capacity.
@@ 5.1. Experimental Setup
Implementation. We develop NeuronMM on top of NeuronX Distributed Inference library (aws-neuron-neuronx-distributed-inference, ) and implement MLP kernels based on NKI (nki2025, ). We evaluate NeuronMM on an trn1.2xlarge instance of Amazon Elastic Compute Cloud (Amazon EC2) equipped with AWS Trainium accelerators, running the Deep Learning Amazon Machine Images (AMI) Neuron (Ubuntu 22.04). We use a single NeuronCore with 16GB HBM on a Trainium chip for evaluation, because using more than one core for inference based on tensor/pipeline/ data parallelism is not fully supported in NKI yet.
Models and datasets. We test LLMs that fit entirely into the HBM and are currently supported by NeuronX Distributed Inference library, including Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B. We evaluate NeuronMM with nine datasets, covering three language modeling datasets (WikiText-2 (merity2016pointer, ), PTB (marcinkiewicz1994building, ), and C4 (raffel2020exploring, )) and six common sense reasoning datasets (OpenBookQA (mihaylov2018can, ), WinoGrande (sakaguchi2021winogrande, ), PIQA (bisk2020piqa, ), HellaSwag (zellers2019hellaswag, ), ARC-e, and ARC-c (clark2018think, )). For fine-tuning with LoRA, we use the yahma/alpaca-cleaned (yahma_alpaca_cleaned, ).
@@ 5.2. Evaluation of $XUV$ Kernel
We compare NeuronMM against two baselines: NKI $XW$ and NKI $XUV$. Both baselines use the state-of-the-art matmul kernel implemented by AWS official (aws_nki_matrix_multiplication, ).
@@ 5.3. Evaluation of MLP Kernel
To assess the trade-off between inference speed and accuracy degradation, we adopt the Speedup Degradation Ratio $\gamma$ (he2024matters, ):
@@ 5.3.1. Impact of Compression Ratio
We evaluate Llama-3.2-1B under compression ratios from 0.05 to 0.5 on the WikiText-2 language modeling task and the ARC Easy commonsense reasoning benchmark. As shown in Figure 6, higher compression ratios reduce the number of model parameters but lead to increased perplexity and decreased accuracy. This demonstrates the fundamental trade-off between model size and predictive performance. The model degradation is acceptable under mild compression. However, as the compression ratio increases performance drops sharply. In particular, between 0.4 and 0.5, PPL surges from 38.87 to 51.29, and accuracy drops from 46% to 40%, indicating that compression has significantly harm the model’s capabilities. To ensure that the accuracy degradation stays within an acceptable range (wang2024svd, ; wang2025svd, ; li2025adasvd, ; yuan2023asvd, ; wang2025dobi, ), we adopt compression ratios of 0.1 and 0.2 for the end-to-end evaluation (Sections 5.3.3-5.3.4).
@@ 5.3.3. Impact of Different LLMs
To demonstrate the generability of NeuronMM across various LLMs, we evaluate it on four LLMs (Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B) across 9 datasets. We report mean accuracy (mAcc) and average end-to-end speedup, with results summarized in Table 4. Across all LLMs evaluated, the SVD-compressed models retain accuracy largely comparable to the original, with mAcc drop $\leq$ 0.10 in every case, a level of loss generally considered acceptable (wang2024svd, ; wang2025svd, ; li2025adasvd, ; yuan2023asvd, ; wang2025dobi, ).
@@ 6. Related Work
SVD and LLM. LLM can have billions of parameters, making inference on resource-constrained hardware challenging. Various model compression techniques have been proposed to reduce their latency and memory footprint (dettmers2023qloraefficientfinetuningquantized, ; Frantar2023OPTQAQ, ; wu2023understandingint4quantizationtransformer, ; frantar2023sparsegpt, ; ma2023llmprunerstructuralpruninglarge, ; guo2023compressostructuredpruningcollaborative, ; ashkboos2024slicegpt, ; DynaBERTHou, ; miles2022informationtheoreticrepresentationdistillation, ; yang2022understandingprivilegedfeaturesdistillation, ). Recent works have explored various SVD-based approaches for model compression. For example, FWSVD (hsu2022language, ) introduces a weighted low-rank factorization scheme, while ASVD (yuan2023asvd, ) develops an activation-aware SVD technique that exploits layer-wise activation patterns to enhance compression effectiveness. SVD-LLM (wang2024svd, ; wang2025svd, ) further incorporates truncation-aware data whitening and layer-specific parameter updates, assigning unique ratios to individual weights. In contrast, Dobi-SVD (wang2025dobi, ) establishes a bijection mapping, allowing the model to automatically learn the best truncation points among layers. However, those works do not consider the characteristics of accelerator hardware and the additional matmul overhead introduced by SVD (e.g., memory spilling) as NeuronMM.
Performance optimization on Trainium. Recent studies (xue2024ninjallmfastscalablecosteffective, ; biran2025adaptiveorchestrationlargescaleinference, ) have demonstrated that AWS Trainium is promising to accelerate generative AI workloads. HLAT (fan2024hlat, ) demonstrates that, through a series of system-level optimizations and training strategies, Trainium can reduce training costs to 60% of those on p4d GPU instances while maintaining comparable model quality to GPU-based baselines. Complementing this, a study (fu2024distributed, ) details NeuronX Distributed Training (NxDT), quantifies scaling and efficiency against contemporary GPU baselines, and elucidates runtime and compiler support critical for stable large-cluster operation.

### v3

@@ 1. Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of text-based tasks (liu2024deepseek; yang2025qwen3). Yet, the steady growth in their parameter counts and architectural complexity makes deployment increasingly prohibitive, particularly in resource-constrained environments. These challenges have driven extensive research into model compression (agarwal2024policy; frantar2023sparsegpt; dettmers2022gpt3; dettmers20218; sanh2020movement; zhang2020accelerating; li2023symbolic; chen2022disco) and efficient hardware design (wan2023efficient; zhu2024survey; jouppi2020domain; bai2023transformer; tzanos2022hardware; tuli2023acceltran; liu2023hardsea; ji2024beta; tuli2023acceltran).
AI accelerators (jouppi2023tpu; coburn2025meta; firoozshahian2023mtia), customized to AI workloads, provide cost-effective and high-performance solutions for training and inference. Trainium is an AI accelerator recently developed by Amazon Web Services (AWS). It has been reported that Trainium can deliver 30–40% lower cost while providing performance comparable to GPU-based EC2 instances (aws_trainium). Each Trainium2 device (the most advanced Trainium) integrates two NeuronCores, each delivering up to 95 TFLOPS of FP16/BF16 compute capability, comparable to NVIDIA A100 GPUs at roughly 60% of the cost. Such a cost–effective advantage makes Trainium an attractive platform for LLM training and inference (fu2024distributed; fan2024hlat). Furthermore, Trainium, as a typical systolic-array architecture, features a programmable memory hierarchy (including two types of on-chip SRAMs and off-chip HBM). It also provides a rich set of specialized compute engines tailored for various AI operators. Such hardware heterogeneity gives programmers a lot of flexibility to explore for better performance.
@@ 2.1. SVD for Weight Compression
SVD is a well-established technique for approximating high-dimensional matrices with low-rank representations (demmel1997applied). Given a weight matrix $W$, SVD factorizes it into three matrices: $U$, $\Sigma$, and $V$, such that $W=U\Sigma V^{\top}$. By retaining only the top-$k$ singular values in $\Sigma$ and their corresponding singular vectors in $U$ and $V$, one obtains a low-rank approximation $W\approx U_{k}\Sigma_{k}V_{k}^{\top}$. This approximation preserves the most informative components of $W$ while substantially reducing the number of parameters to represent $W$. As a result, SVD is particularly well-suited for compressing the large weight matrices in the linear layers of LLMs, where parameter reduction directly improves efficiency with tolerable accuracy loss.
@@ 2.2. AWS Trainium
AWS Trainium is a custom silicon chip designed to accelerate deep learning workloads. It adopts a systolic array–based architecture with rich hardware heterogeneity. Each Trainium chip integrates two NeuronCores. Each NeuronCore functions as an independent heterogeneous compute unit composed of a rich set of specialized engines designed for different operations, such as tensor engine, scalar engine, vector engine, and GPSIMD engine. In addition, Trainium includes DMA engines that can transfer data between HBM and on-chip SRAM in Trainium. Those engines operate in parallel, enabling Trainium to efficiently support diverse deep learning tasks (aws_trainium_arch). In addition, we particularly focus on the tensor engine for matmul in the following discussion.
We use Neuron Kernel Interface (NKI) (aws2024neuron), a bare-metal language and compiler for directly programming NeuronDevices available on AWS Trainium, in our study.
@@
Each tile-level matmul involves two input matrices, named as the stationary (left-hand side) matrix and moving (right-hand side) matrix (using AWS terms). The tensor engine loads the stationary matrix into its internal storage and streams the moving matrix across it. Due to the hardware’s systolic array design, the stationary matrix must be consumed in a transposed layout (aws_trainium_arch). For clarity, we denote the low-level instruction nki.isa.nc_matmul (aws_nki_matrix_multiplication) as NKIMatmul, which carries out this tile-level operation. Thus, computing the product of $AB$ is formulated as NKIMatmul(stationary = $A^{T}$, moving = $B$). The stationary matrix, which remains fixed during the computation, is transposed so that its rows align with the columns of the moving matrix, ensuring both input tiles share the same first dimension which corresponds to the partition dimension in SBUF.
@@
The tile is the fundamental, hardware-native unit of data processed by a single NKIMatmul instruction. The tensor engine imposes strict, hardware-defined maximum dimensions on these tiles, and these constraints differ for the two matrix operands (aws_trainium_arch). Specifically, a tile for the stationary matrix cannot exceed dimensions of (128, 128) for its partition and free axes, respectively. In contrast, a tile for the moving matrix can have a larger free axis, with its dimensions limited to (128, 512). When a matrix is larger than these hardware limits, it must be divided into multiple tiles for processing.
@@ 3.1. Challenge 1: I/O Bottleneck
Executing the SVD-compressed $XUV$ computation with a standard, sequential approach on Trainium results in poor hardware utilization. Figure 4 shows a Neuron Profiler trace of the SVD-compressed up_projection matmul in Deepseek-V3 (liu2024deepseek). For an input length of 4096, with hidden size 7168 and intermediate size 18432, the up_projection matrix ([7168, 18432]) is factorized into two low-rank matrices: $U$ ([7168, 4096]) and $V$ ([4096, 18432]). This transforms the original matmul ([4096, 7168] $\times$ [7168, 18432]) into a chain of three multiplications: [4096, 7168] $\times$ [7168, 4096] $\times$ [4096, 18432]. As highlighted in Figure 4, this decomposition results in frequent idle gaps in Model FLOPs Utilization (MFU) and inefficient use of the on-chip SBUF, indicating that the tensor engine is often stalled waiting for data.
@@ 4. Design
NeuronMM compresses the large weight matrices in MLP layers using block-aligned SVD and restores accuracy through Low-Rank Adaptation (LoRA) (hu2022lora) fine-tuning. It further provides high-performance NKI kernels, termed TrainiumFusion, that execute matmuls on compressed models efficiently by exploiting Trainium’s architectural features.
@@ 4.1. Block-Aligned SVD
Leveraging existing work. Following the standard workflow of post-training LLM compression methods (yuan2023asvd; wang2024svd; ding2025dipsvd), the weight matrix $W$ is first scaled by a matrix $S$ to capture the influence of input activations. $S$ is derived from a random set of input sentences. Specifically, for each MLP layer in LLMs, we record input activations $X$ using forward hooks, compute the covariance matrix, and apply Cholesky decomposition:
@@
where $\text{Model}\_\text{size}_{\text{compressed}}$ and $\text{Model}\_\text{size}_{\text{original}}$ denote the model parameter counts with and without compression. This definition follows the existing work on LLM compression (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; ding2025dipsvd).
@@
Although SVD substantially reduces the size of the weight matrix, it inevitably introduces accuracy loss. To recover accuracy, we apply LoRA fine-tuning to the compressed weights, similar to the prior work (wang2024svd; li2025adasvd). During fine-tuning, we freeze the compressed weights $U_{r}^{\prime}$ and $V_{r}^{\prime}$, and fine-tune them with LoRA:
@@ 4.2.1. XUV NKI Kernel
Using Equations 10-11, we determine the optimal block size. By the roofline model, a kernel becomes compute-bound when its arithmetic intensity exceeds a hardware-specific threshold (williams2009roofline)—for a Trainium NeuronCore with bfloat16 data, this threshold is 222 Flops/Byte (nki_mm2025). Setting the arithmetic intensity (Equation 10) to this threshold yields the minimum block size $B_{M}$ required to saturate the tensor engine. Starting with this $B_{M}$, the peak SBUF usage model (Equation 11) can then identify block combinations that maximize data reuse within the 24 MB SBUF capacity.
@@ 5.1. Experimental Setup
Implementation. We develop NeuronMM on top of NeuronX Distributed Inference library (aws-neuron-neuronx-distributed-inference) and implement MLP kernels based on NKI (nki2025). We evaluate NeuronMM on an trn1.2xlarge instance of Amazon Elastic Compute Cloud (Amazon EC2) equipped with AWS Trainium accelerators, running the Deep Learning Amazon Machine Images (AMI) Neuron (Ubuntu 22.04). We use a single NeuronCore with 16GB HBM on a Trainium chip for evaluation, because using more than one core for inference based on tensor/pipeline/ data parallelism is not fully supported in NKI yet.
Models and datasets. We test LLMs that fit entirely into the HBM and are currently supported by NeuronX Distributed Inference library, including Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B. We evaluate NeuronMM with nine datasets, covering three language modeling datasets (WikiText-2 (merity2016pointer), PTB (marcinkiewicz1994building), and C4 (raffel2020exploring)) and six common sense reasoning datasets (OpenBookQA (mihaylov2018can), WinoGrande (sakaguchi2021winogrande), PIQA (bisk2020piqa), HellaSwag (zellers2019hellaswag), ARC-e, and ARC-c (clark2018think)). For fine-tuning with LoRA, we use the yahma/alpaca-cleaned (yahma_alpaca_cleaned).
@@ 5.2. Evaluation of $XUV$ Kernel
We compare NeuronMM against two baselines: NKI $XW$ and NKI $XUV$. Both baselines use the state-of-the-art matmul kernel implemented by AWS official (aws_nki_matrix_multiplication).
@@ 5.3. Evaluation of MLP Kernel
To assess the trade-off between inference speed and accuracy degradation, we adopt the Speedup Degradation Ratio $\gamma$ (he2024matters):
@@ 5.3.1. Impact of Compression Ratio
We evaluate Llama-3.2-1B under compression ratios from 0.05 to 0.5 on the WikiText-2 language modeling task and the ARC Easy commonsense reasoning benchmark. As shown in Figure 6, higher compression ratios reduce the number of model parameters but lead to increased perplexity and decreased accuracy. This demonstrates the fundamental trade-off between model size and predictive performance. The model degradation is acceptable under mild compression. However, as the compression ratio increases performance drops sharply. In particular, between 0.4 and 0.5, PPL surges from 38.87 to 51.29, and accuracy drops from 46% to 40%, indicating that compression has significantly harm the model’s capabilities. To ensure that the accuracy degradation stays within an acceptable range (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; wang2025dobi), we adopt compression ratios of 0.1 and 0.2 for the end-to-end evaluation (Sections 5.3.3-5.3.4).
@@ 5.3.3. Impact of Different LLMs
To demonstrate the generability of NeuronMM across various LLMs, we evaluate it on four LLMs (Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B) across 9 datasets. We report mean accuracy (mAcc) and average end-to-end speedup, with results summarized in Table 4. Across all LLMs evaluated, the SVD-compressed models retain accuracy largely comparable to the original, with mAcc drop $\leq$ 0.10 in every case, a level of loss generally considered acceptable (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; wang2025dobi).
@@ 6. Related Work
SVD and LLM. LLM can have billions of parameters, making inference on resource-constrained hardware challenging. Various model compression techniques have been proposed to reduce their latency and memory footprint (dettmers2023qloraefficientfinetuningquantized; Frantar2023OPTQAQ; wu2023understandingint4quantizationtransformer; frantar2023sparsegpt; ma2023llmprunerstructuralpruninglarge; guo2023compressostructuredpruningcollaborative; ashkboos2024slicegpt; DynaBERTHou; miles2022informationtheoreticrepresentationdistillation; yang2022understandingprivilegedfeaturesdistillation). Recent works have explored various SVD-based approaches for model compression. For example, FWSVD (hsu2022language) introduces a weighted low-rank factorization scheme, while ASVD (yuan2023asvd) develops an activation-aware SVD technique that exploits layer-wise activation patterns to enhance compression effectiveness. SVD-LLM (wang2024svd; wang2025svd) further incorporates truncation-aware data whitening and layer-specific parameter updates, assigning unique ratios to individual weights. In contrast, Dobi-SVD (wang2025dobi) establishes a bijection mapping, allowing the model to automatically learn the best truncation points among layers. However, those works do not consider the characteristics of accelerator hardware and the additional matmul overhead introduced by SVD (e.g., memory spilling) as NeuronMM.
Performance optimization on Trainium. Recent studies (xue2024ninjallmfastscalablecosteffective; biran2025adaptiveorchestrationlargescaleinference) have demonstrated that AWS Trainium is promising to accelerate generative AI workloads. HLAT (fan2024hlat) demonstrates that, through a series of system-level optimizations and training strategies, Trainium can reduce training costs to 60% of those on p4d GPU instances while maintaining comparable model quality to GPU-based baselines. Complementing this, a study (fu2024distributed) details NeuronX Distributed Training (NxDT), quantifies scaling and efficiency against contemporary GPU baselines, and elucidates runtime and compiler support critical for stable large-cluster operation.